help!!! solve the system by addition -4x+7y=-10
x-5y=9

Answer by checkley71(8403)

-4x+7y=-10
x-5y=9 multiply the second equation by 4 and we get 4x-20y=36 now we add to the first equation thus
-4x+7y=-10
4x-20y=36
--------- which produces
-13y=26 or y=26/-13 or y=-2 & -4x+7(-2)=-10 or -4x-14=-10 or -4x=4 or x=-1
proof -4(-1)+7*(-2)=-10 OR 4-14=-10 OR -10=-10 & -1-5(-2)=9 OR -1+10=9 OR 9=9